VFTOVP(1) VFTOVP(1) NAME vftovp−convertvirtualfont(vf)filestovirtualpropertylists(vpl) SYNOPSIS vftovp[−charcode-format=format][−verbose]vfname[.vf]tfmname[.tfm][ vplname[.vpl]] DESCRIPTION Thismanualpageisnotmeanttobeexhaustive.Thecompletedocumentationforthisversionof TEXcanbefoundintheinfofileormanualWeb2C:ATeXimplementation. Thevftovpprogramtranslatesa(program-oriented)virtualfontfileanditsaccompanyingTEX fontmetricfiletoa(human-oriented)propertylistfile.Thus,thevirtualfontcanbeedited,and itsexactcontentscanbedisplayedmnemonically.ItalsoservesasaVF-filevalidatingprogram; ifnoerrormessagesaregiven,theinputfilesarecorrect. Thefilenamesareextendedwiththeappropriateextensionifnecessary. OPTIONS Theargumentformatto−charcode-formatspecifieshowcharactercodesareoutputintheVPL file.Bydefault,onlylettersanddigitsareoutputusingtheCintegercode(i.e.,inASCII);the othersareoutputinoctal.(Unlessthefont’scodingschemestartswithTeXmathsyorTeX mathex,whenallareoutputinoctal.)Ifformatisascii,allcharactercodesthatcorrespondto graphiccharacters,besidestheleftandrightparentheses,areoutputinASCII.Ontheother hand,ifformatisoctal,allcharactercodesareoutputinoctal. Withoutthe−verboseoption,vftovpoperatessilently.Withit,abannerandprogressreport areprintedonstdout. ENVIRONMENT vftovplooksforvfnameusingtheenvironmentvariableVFFONTS.Ifthatisnotset,ituses thevariableTEXFONTS.Ifthatisnotset,itusesthesystemdefault. Seetex(1)forthedetailsofthesearching. SEEALSO pltotf(1),tftopl(1),vptovf(1). AUTHORS DonaldE.Knuthwrotetheprogram,basedinpartonanideaofDavidFuchs,startingwiththe codefortftopl(1).KarlBerryadapteditforcompilationwithweb2c. Web2C7.32x 19December1994 1